a particle moves along a straight line such that its displacement at any instant 't' is given by x=t3-6t2+3t+4 m.the velocity when acceleration is zero is??

Grade:11

1 Answers

noogler
8 years ago
we know that v= dx/dt , a=dv/dt
v=3t2-12t+3,a=6t-12=0.......t=2
to find v put t=2 in v=3t2-12t+3=-9m/s